Determining Amounts to be Paid on Invoices

Determine the amount to be paid in full settlement of each of the following invoices, assuming that

credit for returns and allowances was received prior to payment and that all invoices were paid within

the discount period:

Merchandise Freight Paid by Seller Returns and Allowance
A. $32,000 FOB destination, n/30 $1,600
B. 12,800 $300 FOB shipping point, 2/10, n/30 2500
C. 21,000 FOB shipping point, 1/10, n/30 4000
D. 9,000 175 FOB shipping point, 2/10, n/30 1000

E. 77,400 FOB destination, 1/10, n/30

A. $

B. $

C. $

D. $

E. $
Answer #1

A. Amount to be paid = $32,000 - $1,600 = $30,400

B. Amount to be paid = ($12,800 - $2,500) – ($10,300 × 2%) + $300 = $10,394

C. Amount to be paid = ($21,000 - $4,000) – ($17,000 × 1%) = $16,830

D. Amount to be paid = ($9,000 – $1,000) – ($8,000 × 2%) + $175 = $8,015

E. Amount to be paid = $77,400 – ($77,400 × 1%) = $76,626